A solid sphere of radius R made of a material of bulk modulus K is surrounded by a liquid in a cylindrical container. A massless pistion of area A floats on the surface of liquid. When a mass M is placed on the piston to compress the liquid the fractional change in the radius of the sphere, d R/R, is…..

  1. A solid sphere of radius R made of a material of bulk modulus K is surrounded by  a liquid in a cylindrical container. A massless pistion of area A floats on the surface of liquid. When a mass M is placed on the piston to compress the liquid the fractional change in the radius of the sphere, d R/R, is…..

Grade:11

1 Answers

Kevin Nash
9 years ago

K = -?P/?V/V

Where ?P = Mg/A ?-?V/V =Mg/AK

? -(Vf-Vi)/Vi = Mg/AK ? Vi-Vf/Vi = Mg/AK

?4/3pR3– 4/3p(R-dR)3/4/3pR3= Mg/AK

?R3- [R3 – 3R2?r]/R3 = Mg/AK

??R/R = Mg/3AK